OpenStudy (anonymous):

\left( sqrt{15x^5y^7} \right)\left( sqrt{5x^3y^5} \right)

OpenStudy (kewlgeek555):

Maybe you should try writing the equation here...

OpenStudy (anonymous):

\(\left( \sqrt{15x^5y^7} \right)\left( \sqrt{5x^3y^5} \right)\)

OpenStudy (anonymous):

\[\left( \sqrt{15x^5y^7} \right)\left( \sqrt{5x^3y^5} \right)\] I know you multiple them across getting \[\sqrt{75x^7y^\left( 13 \right)}\] I dont know where to go from here

OpenStudy (anonymous):

Pull out the perfect square factors

OpenStudy (anonymous):

?

OpenStudy (anonymous):

factors of 75 are 25 and 3. 25 is a perfect square, so... \(\sqrt{75} = 5\sqrt{3}\) Do the same with the variables.

OpenStudy (anonymous):

By the way, you multiplied the variables incorrectly. When multiplying you add the exponents. \(x^5 \times x^3 = x^{5+3} = x^8\)

OpenStudy (anonymous):

ok yes I remeber how to do it with real numbers, what about variables though how can you pull out the squares of variables.

OpenStudy (anonymous):

yes i ment to the power of 8 but the text is small on here so thought the 3 was a 2 is why i got the power of 7, power of 8 is correct

OpenStudy (anonymous):

\(x \times x = x^2\) \(\sqrt{x^2} = x\)

OpenStudy (anonymous):

\(\Large{\left( \sqrt{15x^5y^7} \right)\left( \sqrt{5x^3y^5} \right)}\) `\(\Large{\left( \sqrt{15x^5y^7} \right)\left( \sqrt{5x^3y^5} \right)}\)`

OpenStudy (anonymous):

Any variable squared is a perfect square.

OpenStudy (anonymous):

ok so x would be x to the 4th, how about y since its to the power of 13

OpenStudy (anonymous):

nvm y to the 12th mistype got it thx

OpenStudy (anonymous):

You got it!